OpenStudy (anonymous):

Solving using substitution, medal for help. 2x + 5y = 27 x + y = 12

OpenStudy (anonymous):

Don't give the answer, I'd like to know how to work it out.. :(

OpenStudy (campbell_st):

rewrite the 2nd equation x + y = 12 making y the subject..... then substitute into the 1st equation

OpenStudy (anonymous):

So y = 12 + x ?

OpenStudy (campbell_st):

nearly... to make y the subject you need to subtract x from both sides of the equation. so y = 12 - x does that make sense..?

OpenStudy (anonymous):

Since you moved x to the other side, it has to be the opposite. So it's negative x, right?

OpenStudy (anonymous):

Since I did that, I have 2x + 60 - 5x = 27

OpenStudy (campbell_st):

thats correct... now substitute into the 1st equation 2x + 5(12 - y) = 27 so you need to distribute the 5 as the 1st step then collect like terms... is that ok..?

OpenStudy (campbell_st):

yep... thats perfect... now collect the like terms..

OpenStudy (campbell_st):

oops my mistake its 2x+ 5(12 - x) = 27

OpenStudy (anonymous):

This is what I did: 2x + 5 ( 12 - x ) = 27 2x + 60 - 5x = 27 2x + 5x = 27 + 60 7x = 77 7/7 = 77/7 Leaving me with x = 11

OpenStudy (campbell_st):

not quite.... you need to take care with collecting like terms 2x -5x + 60 = 27 -3x + 60 = 27

OpenStudy (anonymous):

But it's plus 5.. :(

OpenStudy (anonymous):

Never mind, I see where it's -3x..

OpenStudy (campbell_st):

yes it \[5 \times (-x) = -5x\] multiplying a positive and negative results in a negative answer

OpenStudy (anonymous):

so is my x & y value (1,-11) ?

OpenStudy (campbell_st):

ok... so the next step is to do the opposite of + 60 to both sides of the equation... so subtract 60 from both sides... -3x + 60 = 27 -60 -60 --------------------- -3x = -33 now divide both sides by-3 to get the x value...

OpenStudy (anonymous):

It's .. positive 11....

OpenStudy (anonymous):

That was exactly what I got the first time.. :I

OpenStudy (campbell_st):

yep... its more about getting the process right your 1st answer 7x = 27 + 60 which should have been 7x = 87 x = 12 3/7 but remember the process for solving an equation... is do the opposite to find y just substitute your x value into the 2nd equation

OpenStudy (campbell_st):

hope it helped